This question requires coding in RStudio. It is from the book Verzani- Problem 9.9:

The United States Department of Energy conducts weekly phone surveys on the price of gasoline sold in the United States. Suppose one week the sample average was $4.03, the sample standard deviation was $0.42, and the sample size was 800. Perform a one-sided significance test of H0 : \mu = 4.00 against the alternative HA : \mu > 4.00.
Answer #1

Given : Sample size=n=800

Sample mean=\bar{X}=4.03

Sample standard deviation=s=0.42

Hypothesis : H_0:\mu=4.00 Vs H_A:\mu>4.00

The test statistic is ,

Z=\frac{\bar{X}-\mu_0}{s/\sqrt{n}}\to N(0,1)

=\frac{4.03-4.00}{0.42/\sqrt{800}}=2.02

Critical value : Z_{\alpha}=Z_{0.05}=1.64

Decision : Here , Z_{stat}=2.02>Z_{\alpha}=1.64

Therefore , reject the null hypothesis at 0.05 Significance level

Conclusion : Hence , there is sufficient evidence to conclude that the population mean is greater than 4.00.